One day you sewed 2 regular book bags and 4 deluxe book bags and it took you a total of 12.4 hrs. The next day, you sewed 3 regular and 3 deluxe and it took a total of 9.4 hrs. How much time does one regualr bag take, and how much time does one deluxe bag take?

Thank you for your help in solving this question !!!!!!!

Answer by scott8148(6628)

let r=regular time, d=deluxe time ... 2r+4d=12.4 ... 3r+3d=9.4

r+2d=6.2 ... r=6.2-2d ... 3(6.2-2d)+3d=9.4 ... 18.6-6d+3d=9.4 ... -3d=-9.2 ... d=3h4m

3r+3(3h4m)=9h24m ... 3r+9h12m=9h24m ... 3r=12m ... r=4m